OpenStudy (anonymous):

The face of a tent is shaped like a triangle with an area of 11ft2. The area of the triangle is represented by the formula . The height of the tent is 4 feet. How wide is the base of the tent

OpenStudy (anonymous):

so triangles 1/2bh

OpenStudy (anonymous):

1/2(b)(4)=11

OpenStudy (anonymous):

1/2b+2=11 -2 -2 1/2b=11 times 2 (each side) b=22

OpenStudy (anonymous):

Wait.... I made a mistake

OpenStudy (anonymous):

It's 1/2b(4)=11 divide each side by 4 1/2b=2.75 multiply each by 2 b=5.5